*Question 6 GOES WITH QUESTION 5 it’s not a different question *

5. An economy’s AD (Aggregate Demand) function is Y = 1000 – 2P and AS (Aggregate Supply) function is P = 20 + 0.1Y. Show these two lines in a graph. Label graphs. Find the equilibrium price level (P) and GDP (Y) and show them on the graph.


6. Use the AS and AD curves to illustrate your points and discuss the effects of the following events on the price level and on the equilibrium GDP (Y) in the short run:
a) The labor unions negotiate with the government and succeed in increasing the minimum wage.
b) Consumer confidence drops due to an increase in gas-price.
Answer #1

Y= 1000-2P..............Equation 1.

P=20+0.1Y................Equation 2,

P-20

Y=P-20/0.1

1000-2P=P-20/0.1

Solving this equation we get, P= 100

Putting P in any of the equations 1 or 2, we get Equilibrium quantity as 800 units. Refer fig.A. below

6.

Initial equilibrium is at P= 100 and quantity = 800 units (It is Y1 and p1 in diagrams B and C)

a. Increasing minimum wage will cost of production to rise and hence aggregate supply will shift to left. Average price will go up from p1 to p2. and economy will produce less than its potential. This is stagflation as prices are rising and GDP is shifting to left(less than potential). Refer fig. B below.

b. If consumer confidence drops then Aggregate demand will shift to left. Average price levels will come down from P1 to P2 and real GDP will shift to left. This is deflationary pressure in an economy. Refer fig. C below.